OpenStudy (bmk614):

If f is continuous for all real numbers, dy/dx=f(x) and y(2)=4 then y(x)= a) 4+ \[\int\limits_{2}^{x}\] fprime(t) dt b) 4+ \[\int\limits_{2}^{x}\] f(t) dt c) \[\int\limits_{2}^{x}\] f(t) dt-4 d) 4- \[\int\limits_{2}^{x}\] f(t) dt

zepdrix (zepdrix):

Is the first line supposed to contain a capital F somewhere?

OpenStudy (bmk614):

what do you mean?

zepdrix (zepdrix):

All of your options contain F but the information says f, it doesn't make sense.

zepdrix (zepdrix):

Oh you fixed it :P k

OpenStudy (bmk614):

haha yeah sorry about that

zepdrix (zepdrix):

So if \(\large\rm y'=f(x)\) We integrate to get \(\large\rm y=\int f(x)dx+c\) We would then use the data the provided to solve for c, \(\large\rm 4=\int f(2)dx+c\) solving for c gives us, \(\large\rm c=-\int f(2)dx+4\) plugging this back into our general solution gives us,\[\large\rm y=\int\limits f(x)dx+c=\int f(x)dx-\int f(2)dx+4\]

zepdrix (zepdrix):

And I guess we would want to rewrite that as a single integral, using some dummy variable like t,\[\large\rm y(x)=\int\limits_2^x f(t)dt+4\] Mmm sorry to do the whole problem for you, this one was confusing me a lil bit though, needed to write out the steps to see it.

zepdrix (zepdrix):

Any confusion?

OpenStudy (bmk614):

Not at all! Thank you!